OpenStudy (anonymous):

10,000.00 =2000e^0.08t the time needs to be calculate a with continuous compounding how

OpenStudy (curry):

20.18

OpenStudy (dumbcow):

\[2000e^{.08t} = 10,000\] \[e^{.08t} = 5\] \[\ln e^{.08t} = \ln 5\] \[.08t = \ln 5\] \[t = \frac{\ln 5}{.08}\] \[t \approx 20.12\]

OpenStudy (dumbcow):

does that make sense?